# A Watermark for Low-entropy and Unbiased Generation in Large Language Models

### v1

@@ Abstract
Recent advancements in large language models (LLMs) have highlighted the risk of misuse, raising concerns about accurately detecting LLM-generated content. A viable solution for the detection problem is to inject imperceptible identifiers into LLMs, known as watermarks. Previous work demonstrates that unbiased watermarks ensure unforgeability and preserve text quality by maintaining the expectation of the LLM output probability distribution. However, previous unbiased watermarking methods are impractical for local deployment because they rely on accesses to white-box LLMs and input prompts during detection. Moreover, these methods fail to provide statistical guarantees for the type II error of watermark detection. This study proposes the Sampling One Then Accepting (STA-1) method, an unbiased watermark that does not require access to LLMs nor prompts during detection and has statistical guarantees for the type II error. Moreover, we propose a novel tradeoff between watermark strength and text quality in unbiased watermarks. We show that in low-entropy scenarios, unbiased watermarks face a tradeoff between watermark strength and the risk of unsatisfactory outputs. Experimental results on low-entropy and high-entropy datasets demonstrate that STA-1 achieves text quality and watermark strength comparable to existing unbiased watermarks, with a low risk of unsatisfactory outputs. Implementation codes for this study are available online. [^1] https://github.com/djwei96/STA
@@ 1 Introduction
Large language models (LLMs) are large-scale deep learning models that can understand and generate natural languages by learning from a large amount of textual data. Typical generative LLMs, such as ChatGPT [20] and LLaMA [29], can answer questions, translate languages, and create codes with qualities comparable to humans. As LLMs can generate contents more efficiently at a lower cost compared to humans, the risk of LLMs being employed to generate biased, fake, or malicious contents is also increasing [19]. For example, LLMs may exhibit biased information against underrepresented groups of people [2, 7], create misinformation [21], and harm academic integrity [35]. To reduce the harm caused by LLMs, identifying LLM-generated content precisely and efficiently becomes a crucial issue [12].
Watermarks are identifiers imperceptible to humans but detectable by certain models [18]. In particular, in the era of LLMs, a watermarking method is a strategy to alter the logits or probability distribution during token generation by LLMs [12, 8], with the randomness controlled confidentially by LLM owners. Statistical hypotheses are constructed with the knowledge advantage for the randomness (as the alternative hypothesis) and without the knowledge advantage (as the null hypothesis). For example, previous work (denoted as KGW) [12] first divides the token set into a green and a red list at each generation step. Next, it increases the logits output by the pretrained LLM for green list tokens, leading the watermarked LLM to generate tokens in the green list with higher probabilities. The knowledge advantage in KGW is the randomness of the green-red-list partition at each step.
@@
In practice, watermarks should have unforgeability against deciphering watermarking generation attacks [18]. A watermarking method demonstrates the ability against forgeries if it can hide the distinguishability between the original unwatermarked text and its watermarked counterpart [5, 18]. Thus, it is required that a watermarking method adjusts the probability distribution while maintaining the same expectation as the unwatermarked distribution [11, 15], defined as unbiased watermarks.
Hu et al. [11] find KGW is biased under the expectation maintaining requirement. In response, they propose an unbiased $\gamma$-reweight method that alters the probability distribution at each generation step. To detect watermarked texts, they use the log likelihood ratio (LLR) test by comparing the likelihood of the text produced by watermarked and unwatermarked white-box LLMs. Their knowledge advantage to construct statistical tests is the controlled randomness in $\gamma$-reweight. Wu et al. [33] propose Dipmark, an extension of $\gamma$-reweight with more general parameter settings.
Existing unbiased watermarking methods focus on the unbiased reweighting of the initial probability distribution [11, 33]. However, these methods face challenges for the following reasons. First, to accurately compute the likelihood ratio, methods employing the LLR test require the prompt as input and a white-box LLM even in watermark detection [8, 11]. The requirement of a white-box LLM is costly and impractical for local deployment. Also, relying on the LLR test makes watermarks unstable because changing the first token of the generated text can lead to huge deviations from the original likelihood value, due to the autoregressive mechanism of LLM inference [8]. Second, although both $\gamma$-reweight and Dipmark ensure the type I error of watermark detection, they fail to provide statistical guarantees for the type II error [11, 33]. That is, they can only guarantee that unwatermarked texts are not misclassified, but cannot guarantee that all watermarked texts are detectable.
@@
In response to the above challenges, we propose the Sampling One Then Accepting (STA-1) method for watermarking LLMs. STA-1 traces back to the original KGW method where the token set is divided into a green and a red list at each generation step [12]. Instead of raising logits in the green list, STA-1 samples a token from the original probability distribution and accepts it if it is in the green list. If the sampled token is in the red list, it resamples another token and accept it. By counting the number of green list tokens, it employs the $z$-test for watermark detection, which naturally addresses the robustness issue encountered with the LLR test and eliminates the need for prompts and white-box LLMs in detection [12]. Meanwhile, we prove that STA-1 is unbiased and provides implicit statistical guarantees for the type II error. More interestingly, the bounds are linked to the Gini index of the probability distribution, which is a common metric in machine learning [3] compared to the proposed Spike entropy in previous work [12].
@@
In this study, we also clarify the watermark strength and text quality tradeoff in unbiased watermarks. The KGW method faces a tradeoff between watermark strength and text quality, which means a higher detection power results in a lower text quality [12]. Previous work claims that unbiased watermarks can avoid this tradeoff given the preserved text quality by maintaining the expectation of probability distribution [11]. We challenge this claim by considering a simple low-entropy scenario, where we show that unbiased watermarks still face a tradeoff between watermark strength and text quality. However, the text quality is now related to the risk of unsatisfactory outputs. Specifically, unsatisfactory outputs in low-entropy scenarios represent that the watermarking method alters probability distribution such that high-probability tokens cannot be sampled at risk. We discuss the risk via the variance of the probability after altering, which is a common practice of risk-return analysis [27]. We prove that STA-1 is less risky than previous unbiased watermarks. Moreover, we propose STA-M, an extension of STA-1, by setting up a threshold for entropy in generation [16, 30] and sampling more times for high-entropy steps. Although STA-M is not unbiased theoretically, it allows higher watermark strength with small performance shifts empirically. Also, STA-M is robust against various watermarking attacks. We summarize our contributions as follows:
@@
  • We propose STA-1, an unbiased watermarking method that is practical and has statistical guarantees on type II error of watermark detection. Moreover, we introduce STA-M, an extension of STA-1 that enhances watermark strength with low text quality shifts.
@@
  • We clarify the watermark strength and text quality tradeoff in unbiased watermarks. In low-entropy scenarios, the text quality is related to the risk of unsatisfactory outputs. We show that STA-1 has a lower risk theoretically compared to other unbiased watermarks.
@@
  • Experimental results on public low-entropy and high-entropy datasets empirically show that STA-1 achieves comparable performances against other unbiased watermarks and has a low risk of unsatisfactory outputs. Meanwhile, STA-M demonstrates high watermark strength in the low-entropy dataset and robustness against different watermarking attacks.
@@ 2 Preliminary
Notations. We follow notations in previous work [12, 11] to represent the generation task of LLMs. Let $P_{M}$ denote a pretrained LLM and $\mathcal{V}$ is the overall token (vocabulary) set. An example token set contains more than 50,000 tokens ($|\mathcal{V}|>50000$) [23]. For simplicity, we use Python-style notation for an ordered token sequence, where $x^{-m:n}=(x^{-m},x^{-m+1},\cdots,x^{n})$, $m$ and $n$ are integers. In a typical LLM generation task, an LLM receives a sequence of $N_{p}+1$ tokens $x^{-N_{p}:0}$, known as a prompt, and outputs a sequence of $T$ tokens $x^{1:T}$ step by step. At step $t$, the probability of each token in the token set $\mathcal{V}$ is given by the conditional distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$. The LLM generation follows an autoregressive fashion, where the joint probability of the generated tokens are as follows
$P_{M}(x^{1:T}|x^{-N_{p}:(-1)})=\prod_{t=1}^{T}P_{M}(x^{t}|x^{-N_{p}:(t-1)}).$ (1)
When applying watermarking techniques, the LLM employs a private key $k$ to adjust the conditional distribution from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $P_{M,w}$ indicates a watermarked model and the private key $k$ is randomly selected from a key space $K$ according to a known distribution $P_{K}(k)$. An unbiased watermark requires that the expectation of the watermarked distribution equals that of the original distribution [11, 18], defined as follows.
@@
Previous distribution reweighting methods. Many existing watermarking methods fall into the category of distribution reweighting [11, 12, 16, 33]. Formally, a reweighting function $R_{k}:\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ maps from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $\mathcal{P}_{\mathcal{V}}$ denotes the probability distribution space over the vocabulary set $\mathcal{V}$. A reweighting method $R:K\times\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ contains all realized reweighting functions $R_{k}$ among the key space $k\in K$. Following Definition 1, $R$ is an unbiased reweighting method if $\mathbb{E}_{k\sim P_{K}(k)}\left[R_{k}(P_{M})\right]=P_{M}$. Next, we introduce previous distribution reweighting methods [11, 12, 33] and refer readers to Appendix A for more details.
KGW [12]: The KGW method [12] randomly splits the vocabulary set $\mathcal{V}$ into a green list and a red list based on a uniformly distributed key $k$. The soft KGW method adds a predefined constant $\delta$ to the green list tokens’ logits while keeping the red list tokens’ logits fixed.
Dipmark [33] and $\gamma$-reweight [11]: Wu et al. [33] propose an unbiased watermarking method named Dipmark. Dipmark shuffles all probability masses $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ over the vocabulary set within the probability interval $[0,1]$ based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probabilities in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Dipmark becomes $\gamma$-reweight when $\alpha=0.5$.
@@ 3 A Simple Protocol for a Low-entropy Scenario
The low-entropy text refers to a relatively deterministic sequence in natural language. The entropy measures the uncertainty of the probability distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ at a single generation step among the token set $\mathcal{V}$, where low entropy means low uncertainty. For example, in code writing, the structure of a code sequence is regularized where few changes can be made [16]. More explicitly, for a typical English pangram such as ‘The quick brown fox jumps over the lazy dog’ [12], both humans and machines should generate similar if not identical output. For example, when provided with the prompt ‘The quick brown fox jumps over the lazy’, the trained LLaMA-2-7B [29] outputs an empirical probability above 0.8 for the next word ‘dog’.
Figure 1: Example of a low-entropy scenario.
Problem modeling. Low-entropy scenarios exist in text generation tasks of LLMs. We aim to model a simple problem protocol for the low-entropy generation scenario. For simplicity, we consider the low-entropy scenario where only one token probability is significantly large. Specifically, denote $p_{max}$ as the largest probability of a token in the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. We make an intuitive assumption that except $p_{max}$, other $|\mathcal{V}|-1$ probabilities are small enough to uniformly fill in the remaining $1-p_{max}$ probability value. Figure 1 demonstrates the low-entropy scenario given the prompt ‘The quick brown fox jumps over the lazy’.

### v2

@@ Abstract
Recent advancements in large language models (LLMs) have highlighted the risk of misusing them, raising the need for accurate detection of LLM-generated content. In response, a viable solution is to inject imperceptible identifiers into LLMs, known as watermarks. Previous work demonstrates that unbiased watermarks ensure unforgeability and preserve text quality by maintaining the expectation of the LLM output probability distribution. However, previous unbiased watermarking methods suffer from one or more of the following issues: (1) requiring access to white-box LLMs during detection, (2) incurring long detection time, (3) being not robust against simple watermarking attacks, (4) failing to provide statistical guarantees for the type II error of watermark detection, and (5) being not statistically unbiased for low-entropy scenarios, which hinder their deployment in practice. This study proposes the Sampling One Then Accepting (STA-1) method, a watermark that can address all of these issues. Moreover, we discuss the tradeoff between watermark strength and text quality for unbiased watermarks. We show that in low-entropy scenarios, unbiased watermarks face a tradeoff between watermark strength and the risk of unsatisfactory outputs. Experimental results on both low-entropy and high-entropy datasets demonstrate that STA-1 achieves text quality and watermark strength comparable to existing unbiased watermarks, with a low risk of unsatisfactory outputs. Implementation codes for this study are available online. [^1] https://github.com/djwei96/STA
@@ 1 Introduction
Large language models (LLMs) are large-scale deep learning models that can understand and generate natural languages by learning from a large amount of textual data. Typical generative LLMs, such as ChatGPT (Ouyang et al., 2022) and LLaMA (Touvron et al., 2023), can answer questions, translate languages, and create codes with qualities comparable to humans. As LLMs can generate contents more efficiently at a lower cost compared to humans, the risk of LLMs being employed to generate biased, fake, or malicious contents is also increasing (Mirsky et al., 2023). For example, LLMs may exhibit biased information against underrepresented groups of people (Abid et al., 2021; Fang et al., 2024), create misinformation (Pan et al., 2023), and harm academic integrity (Zhao et al., 2023). To reduce the harm caused by LLMs, identifying LLM-generated content precisely and efficiently becomes a crucial issue (Kirchenbauer et al., 2023a).
@@
Watermarks are identifiers imperceptible to humans but detectable by certain models (Liu et al., 2023b). In the era of LLMs, a watermarking method is a strategy to control the randomness of token generation by LLMs (Kirchenbauer et al., 2023a; Fernandez et al., 2023), with the randomness preserved confidentially by LLM owners. In practice, watermarks should have unforgeability against deciphering watermarking generation attacks (Liu et al., 2023b). A watermarking method demonstrates the ability against forgeries if it can hide the distinguishability between the original unwatermarked text and its watermarked counterpart (Christ et al., 2023; Liu et al., 2023b). Thus, it is required that a watermarking method adjusts the probability distribution while maintaining the same expectation as the unwatermarked distribution (Hu et al., 2024; Kuditipudi et al., 2023), defined as unbiased watermarks.
Existing unbiased watermarks can be categorized according to the stage where watermarks are injected: distribution reweighting and controlled sampling (Liu et al., 2023b). For distribution reweighting, Hu et al. (2024) proposes $\gamma$-reweight, which uses the log-likelihood ratio (LLR) test by comparing the likelihood of the text produced by watermarked and unwatermarked white-box LLMs. It requires the prompt as input and a white-box LLM in watermark detection (Fernandez et al., 2023; Hu et al., 2024). Also, the watermark is unstable because changing the first token of the generated text can lead to huge deviations from the original likelihood value (Fernandez et al., 2023). In response, Wu et al. (2024) avoid the LLR test and propose Dipmark, an extension of $\gamma$-reweight with more general parameter settings. However, although both $\gamma$-reweight and Dipmark ensure the type I error of watermark detection, they fail to provide statistical guarantees for the type II error (Hu et al., 2024; Wu et al., 2024). For controlled sampling, Christ et al. (2023) introduce a watermarking method that uses a sequence of random values to guide the token sampling process. However, their method is not robust enough against simple removal attacks (Liu et al., 2023b). Kuditipudi et al. (2023) also use random values to control the sampling and introduce a permutation test on detection that does not require white-box access to LLMs. However, this permutation test is time-consuming theoretically and empirically. Fairoze et al. (2023) propose to sample the token sequence generation until its hash matches a key value. According to their distortion-free definition, the upper bound of the difference between probabilities before and after watermarking is $\exp(-a)$, where $a$ is the minimal entropy. The difference is not negligible in low-entropy scenarios. Note that using random values to control sampling can be treated as a special case of distribution reweighting where only the probability of the sampled token is reweighted to 1 (Kuditipudi et al., 2023). Thus, we build our analysis framework in Section 2 solely based on distribution reweighting.
@@
We conclude the research gaps in Table 5 in Appendix A. In response to the challenges, we propose the Sampling One Then Accepting (STA-1) method that can simultaneously overcome the above gaps. STA-1 traces back to the original watermarking method (denoted as KGW) (Kirchenbauer et al., 2023a) where the token set is divided into a green and a red list at each generation step. Instead of raising logits in the green list, STA-1 samples a token from the original probability distribution and accepts it if it is in the green list. If the sampled token is in the red list, it resamples another token and accept it. By counting the number of green list tokens, it employs the $z$-test for watermark detection, which naturally addresses the issues encountered with the LLR test and eliminates the need for prompts and white-box LLMs in detection. Simultaneously, the detection of STA-1 is efficient and only requires $O(m)$ time complexity, where $m$ is the number of tokens. The STA-1 method is also robust against simple insertion and removal attacks because changing tokens can only affect the detection score around these tokens. Meanwhile, we prove that STA-1 is unbiased and provides statistical guarantees for the type II error. More interestingly, the bounds are linked to the Gini index of the probability distribution, which is a common metric in machine learning (Breiman, 2017) compared to the proposed Spike entropy in previous work (Kirchenbauer et al., 2023a).
@@
In this study, we also clarify the watermark strength and text quality tradeoff in unbiased watermarks. The KGW method faces a tradeoff between watermark strength and text quality, which means a higher detection power results in a lower text quality (Kirchenbauer et al., 2023a). Previous work claims that unbiased watermarks can avoid this tradeoff given the preserved text quality by maintaining the expectation of probability distribution (Hu et al., 2024). We challenge this claim by considering a simple low-entropy scenario, where we show that unbiased watermarks still face a tradeoff between watermark strength and text quality. However, under the same expectation constraint, the text quality is related to the risk of unsatisfactory outputs. Specifically, unsatisfactory outputs in low-entropy scenarios represent that the watermarking method alters probability distribution too much such that high-probability tokens cannot be sampled at risk. We discuss the risk via the variance of the probability after altering, which is a common practice of risk-return analysis (Sharpe, 1998). We prove that STA-1 is less risky than previous unbiased watermarks. Moreover, we propose STA-M, an extension of STA-1, by setting up a threshold for entropy in generation (Lee et al., 2023; Wang et al., 2023) and sampling more times for high-entropy steps. Although STA-M is not unbiased theoretically, it allows higher watermark strength with small performance shifts empirically. Also, STA-M is robust against various watermarking attacks. We summarize our contributions as follows:
@@
1. We propose STA-1, an unbiased watermarking method that is practical and has statistical guarantees on type II error of watermark detection. Moreover, we introduce STA-M, an extension of STA-1 that enhances watermark strength with low text quality shifts.
@@
2. We clarify the watermark strength and text quality tradeoff in unbiased watermarks. In low-entropy scenarios, the text quality is related to the risk of unsatisfactory outputs. We show that STA-1 has a lower risk theoretically compared to other unbiased watermarks.
@@
3. Experimental results on public low-entropy and high-entropy datasets empirically show that STA-1 achieves comparable performances against other unbiased watermarks and has a low risk of unsatisfactory outputs. Meanwhile, STA-M demonstrates high watermark strength in the low-entropy dataset and robustness against different watermarking attacks.
@@ 2 Preliminary
Notations. We follow notations in previous work (Kirchenbauer et al., 2023a; Hu et al., 2024) to represent the generation task of LLMs. Let $P_{M}$ denote a pretrained LLM and $\mathcal{V}$ is the overall token (vocabulary) set. An example token set contains more than 50,000 tokens ($|\mathcal{V}|>50000$) (Radford et al., 2019). For simplicity, we use Python-style notation for an ordered token sequence, where $x^{-m:n}=(x^{-m},x^{-m+1},\cdots,x^{n})$, $m$ and $n$ are integers. In a typical LLM generation task, an LLM receives a sequence of $N_{p}+1$ tokens $x^{-N_{p}:0}$, known as a prompt, and outputs a sequence of $T$ tokens $x^{1:T}$ step by step. At step $t$, the probability of each token in the token set $\mathcal{V}$ is given by the conditional distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$. The LLM generation follows an autoregressive fashion, where the joint probability of the generated tokens are as $P_{M}(x^{1:T}|x^{-N_{p}:0})=\prod_{t=1}^{T}P_{M}(x^{t}|x^{-N_{p}:(t-1)})$.
When applying watermarking techniques, the LLM employs a private key $k$ to adjust the conditional distribution from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $P_{M,w}$ indicates a watermarked model and the private key $k$ is randomly selected from a key space $K$ according to a known distribution $P_{K}(k)$. An unbiased watermark requires that the expectation of the watermarked distribution equals that of the original distribution (Hu et al., 2024), defined as follows.
@@
Previous distribution reweighting methods. Since controlled sampling can be viewed as a special case of distribution reweighting, we build our analysis framework based on distribution reweighting. Formally, a reweighting function $R_{k}:\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ maps from $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ to $P_{M,w}(x^{t}|x^{-N_{p}:(t-1)};k)$, where $\mathcal{P}_{\mathcal{V}}$ denotes the probability distribution space over the vocabulary set $\mathcal{V}$. A reweighting method $R:K\times\mathcal{P}_{\mathcal{V}}\rightarrow\mathcal{P}_{\mathcal{V}}$ contains all realized reweighting functions $R_{k}$ among the key space $k\in K$. Following Definition 1, $R$ is an unbiased reweighting method if $\mathbb{E}_{k\sim P_{K}(k)}\left[R_{k}(P_{M})\right]=P_{M}$. Next, we introduce previous watermarking methods (Kirchenbauer et al., 2023a; Hu et al., 2024; Wu et al., 2024; Kuditipudi et al., 2023) and refer readers to Appendix B for more details.
KGW (Kirchenbauer et al., 2023a): The KGW method (Kirchenbauer et al., 2023a) randomly splits the vocabulary set $\mathcal{V}$ into a green list and a red list based on a uniformly distributed key $k$. The soft KGW method adds a predefined constant $\delta$ to the green list tokens’ logits while keeping the red list tokens’ logits fixed.
Dipmark (Wu et al., 2024) and $\gamma$-reweight (Hu et al., 2024): Wu et al. (2024) propose an unbiased watermarking method named Dipmark. Dipmark shuffles all probability masses $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ over the vocabulary set within the probability interval $[0,1]$ based on a key $k$. A hyperparameter $\alpha\in[0,0.5]$ partitions the interval $[0,1]$ into three segments: $[0,\alpha]$, $(\alpha,1-\alpha]$, and $(1-\alpha,1]$. Probabilities in the first segment are set to 0, those in the second remain constant, and those in the third are doubled. Dipmark becomes $\gamma$-reweight when $\alpha=0.5$.
RDW (Kuditipudi et al., 2023): We focus on the RDW method via an inverse transform sampling scheme. Given a uniformly distributed key $k$, RDW first shuffles all probability masses within the interval $[0,1]$, then it randomly samples a value $u\sim\text{U}(0,1)$. Here, $u$ is viewed as the cumulative distribution function value of $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ with respect to the permutation; it is subsequently inverse transformed to generate a token. The probability of the sampled token is reweighted to 1, while the probabilities of all other tokens are reweighted to 0.
@@ 3 A Simple Protocol for a Low-entropy Scenario
The low-entropy text refers to a relatively deterministic sequence in natural language. The entropy measures the uncertainty of the probability distribution $P_{M}(x^{t}|x^{-N_{p}:(t-1)})$ at a single generation step among the token set $\mathcal{V}$, where low entropy means low uncertainty. For example, in code writing, the structure of a code sequence is regularized where few changes can be made (Lee et al., 2023). More explicitly, for a typical English pangram such as ‘The quick brown fox jumps over the lazy dog’ (Kirchenbauer et al., 2023a), both humans and machines should generate similar if not identical output. For example, when provided with the prompt ‘The quick brown fox jumps over the lazy’, the trained LLaMA-2-7B (Touvron et al., 2023) outputs an empirical probability above 0.8 for the next word ‘dog’.
Problem modeling. Low-entropy scenarios exist in text generation tasks of LLMs. We aim to model a simple problem protocol for the low-entropy generation scenario. For simplicity, we consider the low-entropy scenario where only one token probability is significantly large. Specifically, denote $p_{max}$ as the largest probability of a token in the probability distribution $P_{M}(\cdot|x^{-N_{p}:(t-1)})$. We make an intuitive assumption that except $p_{max}$, other $|\mathcal{V}|-1$ probabilities are small enough to uniformly fill in the remaining $1-p_{max}$ probability value.